Hobart 26m

Module: HOB+1004, start pos: ???? GB

Major antenna failure at start of experiment, not participating at present (Ross)

  • 04:15 UT: Hobart rejoined schedule at 0415 after repairs made to emergency stop system. No useful data recorded prior to this time. (JMc)
    • Disk position at start of 0415 scan was 1767 TB (JMc)
  • 06:58 UT: First module (USN-0111) ran out of space, FS automatically swapped to HOB+1004. Starting at 6377 GB. (Dave)
  • 07:00 UT: 20K stage at 40K. (Dave)
  • 08:00 UT: Cryo temp continuing to rise. Driving to Mt.P to turn on vacuum pump. (Dave)
  • 09:00 UT: Cryo pressure now good. Temps going down. (Dave)


Module: ???+????, start pos: 2417.620 GB

21:27 UT: large clock jump (Ross).

Module: HOB+0029


Module: USN+0224, start pos: 2298.335 GB